OpenStudy (jymoore24):

Sandy Owns a restaurant. She Charges $5.50 for 3 eggs and 2 pieces of toast and $3.50 for 2 eggs and 1 piece of toast A. How much does Sandy Charge per egg? B. How Much does Sandy Charge per piece of toast? Need Help Will Fan And Medal

OpenStudy (toxicwaste):

$1.83 per egg

OpenStudy (xguardians):

$5.50 = 3e + 2t $3.50 = 2e + 1t Solve for t in second. $3.50 - 2e = 1t Subsitute $5.50 = 3e + 2($3.50-2e) $5.50 = 3e + $7.00 -4e -$1.50 = -e e = $1.50 Resubstitute $5.50 = 3($1.50) + 2t $5.50 = $4.50 + 2t $1.00 = 2t t = $0.50 Eggs are $1.50 each, toast is $0.50 each.

OpenStudy (jymoore24):

Wow thank you so much that helped alot @xGuardians